OpenStudy (anonymous):

write an explicit formula for the sequence 7,2,-3,-8,-13,... Then find a^14

OpenStudy (michele_laino):

note that, your sequence is an arithmetic progression, whose first term is 7, and constant c=-5, so generic term a_n, is given by the subsequent formula: \[a _{n}=a _{1}+(n-1)*c=7+(n-1)*(-5)=\] \[=7-5n+5=12-5n\] so, please, substitute n=14, and you will get your answer

OpenStudy (anonymous):

ok i got 98

OpenStudy (michele_laino):

no, sorry: \[a _{14}=12-5*14=...\]

OpenStudy (michele_laino):

\[a _{14}=-58\] @bandnerd24

OpenStudy (anonymous):

oh well thanks for the help

OpenStudy (michele_laino):

thanks!